# GRAMMAR: Grounded and Modular Methodology for Assessment of Domain-Specific Retrieval-Augmented Language Models

### v4

@@ 3.2 From Templates To Evaluation Data
Figure 2: Scaled Generation of Query-answer Pairs With Combinatorial Power. The step 3 and step 4 are independent of each other and depend only on the SQL templates.
@@ 3.3 Modular Evaluation
Assessing an RAG model’s overall performance may obscure insights into the robustness or weaknesses of its constituent modules (refer to Section 4.2). It is essential, therefore, to evaluate the contributions of the retrieval and language modeling (LM) components individually.
We introduce three group types to distinguish the knowledge gap and the robustness issue, as shown in Figure 1. All textual queries $Q$ generated from a specific SQL logic $q_{\mathrm{sql}}$ can be grouped semantically. [^1] Refer to Appendix D for an illustration. Depending on the model $M$’s performance, we categorize $q_{\mathrm{sql}}$ (and consequently, the group $Q$) into three principal categories:
@@
    $\forall q_{t}\in S,\neg M\left(q_{t}\right)$
    wherein $\neg M\left(q_{t}\right)$ signifies an inaccurate response from $M$ to the text query $q_{t}$.
@@
  • Non-robust Groups are characterized by model $M$’s exhibited capability to potentially address the query logic, yet there exists at least one query $q_{t}$ within $S$ that $M$ inaccurately predicts. This is formally denoted as:
    $\left(\exists q_{t}\in S,M(q_{t})\right)\land\left(\exists q_{t}\in S,\neg M(q_{t})\right)$
Defective Modules Responsible for The Robustness Issue
Gap groups mostly reflect the limited coverage of the retrieval database. The analysis of the semantics of queries in gap groups can help expand the database. However, defective modules responsible for the robustness issue need to be further identified by answering the following questions: Does the retrieval module source and provide the sufficient context needed for a given query? With the sufficient context at its disposal, is the LM capable of drawing accurate inference? Utilizing a “context comparison” methodology enables pinpointing the LM’s role in generating erroneous predictions when adequate context is provided. By contrasting the document indices $\text{idx}_{c}$ of each non-robust instance against indices from accurate predictions within analogous semantic groups against those linked to correct predictions within the same semantic groups $I=\text{idx}_{1},\text{idx}_{2},\dots$, we evaluate the sufficiency of context. This assumes the LM’s reliance on external information for query responses, mirroring our scenario and prevalent industrial applications. Should any index in $I$ match $\text{idx}_{c}$, we reclassify the outcome as a successful retrieval yet an LM shortfall. This strategy allows for precise identification of LM-related non-robust instances. Nevertheless, the absence of such matches does not conclusively indicate retrieval failures. Given our experimental setup with a synthetic database comprising 10 documents, this “context comparison” method effectively identifies all document combinations across a moderate query range (3 to 6) within semantically related groups. This framework lays the groundwork for future explorations into more intricate cases.
A reliable robustness metric is necessary to allow precise identification of detective modules. We quantify robustness as the fraction of instances correctly answered, excluding those in gap groups:
@@
$\begin{split}R=\frac{\text{\# with Correct Predictions}}{\text{Total \#}-\text{Total \# in Gap Groups}}\end{split}$ (1)
, where “#” represents “the number of instances”. By omitting gap groups, this approach ensures that the metric reflects true robustness, distinct from knowledge gaps. Consequently, accuracy Acc is articulated as:
$\begin{split}\text{Acc}=R\times\frac{\text{Total \#}-\text{Total \# in Gap Groups}}{\text{Total \#}}\end{split}$ (2)
This formulation reveals how the proportion of instances within gap groups $\lambda$ influences the efficacy of accuracy as a measure of robustness. Specifically, a higher $\lambda$ value in the robust compared to the non-robust cluster may obscure the model’s non-robust characteristics. Section 4 will empirically demonstrate the crucial importance of excluding gap groups for accurately assessing system vulnerabilities.
4 Case Study: Question Answering on Industrial Projects
This section presents a case study focusing on question answering about industrial projects. It serves to critically evaluate both existing reference-free evaluation frameworks and the newly proposed evaluation framework
System Overview
Our study incorporates two distinct systems: 1) Dense Retrieval + GPT3.5  [^2] OpenAI’s gpt-3.5-turbo-16k: It is deployed in a commercial context with secure access to the retrieval module via the Qdrant API, ensuring data privacy. The system employs an OpenAI embedding model for document embedding, providing a basis to assess the reference-free evaluation framework’s reliability (§4.1). 2) Simple Sparse Retrieval + GPT3.5: The absence of comprehensive research into the assured robustness of dense retrieval systems makes it difficult to pinpoint distinct robust or vulnerable characteristics with confidence. Consequently, we employ a straightforward keyword-matching retrieval approach. Although this method is less prevalent in real-world applications, this method’s inherent shortfall in handling lengthy queries makes it exceptionally ideal to validate our evaluation framework in conducting reliable hypothesis testing (§4.2).
Synthetic Knowledge Base
Concurrently, a synthetic dataset has been created, consisting of project documents from the fictitious company, Aurp, serving as a retrieval database. It facilitates a precise evaluation of the retrieval component and the language model, since the knowledge contained within the synthetic data is not present in the training datasets of large language models. In scenarios where retrieval fails to procure adequate context, the language models are definitely unable to accurately respond to queries. Additional information on this methodology can be found in Appendix F.
Mapping Domain Knowledge into Database Schema
Our approach distills domain knowledge into a database schema centered around three pivotal entity types in project management: Project, Client, and Employee.  [^3] This triad are reflective of the core components outlined in the Project Management Body of Knowledge (PMBOK), a widely recognized standard in the field of project management (Guide, 2001).This schema encapsulates the quintessential entity types found within a corporate setting. With the scalability of our framework, the limited amounts of entity types and associated attributes and entities lead to large amounts of instances for evaluation: 198 examples for the commercial case (with totally 12 attributes and 11 rows) and 942 examples for the synthetic scenario (with 14 attributes and 49 rows in total). The details of the database schema and the amounts of columns and rows for each table are shown in Appendix E and G), respectively.
4.1 Is Reference-free Evaluation Reliable?
Leveraging ground truths, our evaluation framework sets a benchmark for reference-free assessments. This section delves into the reliability of reference-free evaluation methods, utilizing examples generated for the commercial case. The focus is to discern the reliability of such evaluation approaches: How trustworthy are these reference-free protocols? From what perspectives can they be relied upon in domain-specific contexts?
Reference-free Evaluation Protocols
We assess two reference-free evaluation protocols: SelfCheck and RAGAS. RAGAS-Fact utilizes the context-query-response triplets to assess the veracity of responses. It evaluates the faithfulness of a response by calculating the ratio of claims grounded on the context to the total claims made. This process, termed as Ragas-Fact, involves identifying statements that hold atomic facts, following the methodology outlined by Chern et al. (2023). SelfCheck, as introduced by Manakul et al. (2023), relies on stochastically generating responses from the model under the premise that incorrect answers are unlikely to repeatedly yield identical outcomes through random sampling. This principle, initially applied to LLMs, is adapted for our analysis of RAG systems. In a departure from its original application, we enhance the evaluation prompt to include not only stochastic responses but also the query itself. Refer to Appendix H for details. To generate four stochastic samples, we adjust the temperature setting to 1.0, contrasting with a temperature of 0.0 used to procure the primary response. A response is deemed correct if it aligns consistently across all stochastic samples.
RAGAS’ Reliable Assessment of Correct Predictions Amidst Broader Unreliability in Reference-Free Evaluation
To gauge reliability, we devise two metrics: the ratio of predictions judged as accurate among all those classified as correct by the evaluation, and the ratio of predictions deemed accurate within the set of instances that were indeed correctly predicted. When considering reference-free evaluation as a binary classification task, these metrics align with precision and recall, respectively. In this context, the positive class represents correct predictions, and the negative class represents incorrect ones. The ground truth corresponds to the actual accuracy of model predictions, while the evaluation task is to determine the correctness of these predictions. As shown in Table 1, RAGAS demonstrates effectiveness in assessing accurately predicted instances, yet it achieves only a 19% accuracy rate across all evaluations, highlighting a deficiency in correctly identifying erroneous predictions. This trend suggests a propensity for inflating model performance assessments, likely due to the RAG system producing responses that, while contextually relevant, fail to directly address the intended query. SelfCheck performs even worse than RAGAS in both cases. SelfCheck’s performance falls short of RAGAS’s, underperforming in both precision and recall. The preliminary insight suggests that reference-free evaluation methods are inherently prone to inaccuracies, underscoring the critical need for a robust, reference-based evaluation framework to ensure dependable model assessment.
Precision Recall
Ours (Upper Bound) 1 1
RAGAS-Fact 19% 92%
SelfCheck 15% 50%
Table 1: Reliability of reference-free evaluation protocols.
Robust Non-robust Success?
Accuracy Baseline: Unbalanced Sampling (US) 0.31 0.43 No
US+Modular Evaluation (Retrieval) 0.36 0.46 No
Stratified Sampling (SS) 0.31 0.29 Yes
SS+Modular Evaluation (Retrieval) 0.38 0.35 Yes
Robustness Baseline: Unbalanced Sampling (US) 0.84 0.85 No
US+Modular Evaluation (Retrieval) 0.99 0.93 Yes
Stratified Sampling (SS) 0.78 0.74 Yes
SS+Modular Evaluation (Retrieval) 0.94 0.87 Yes
Table 2: Evaluating the efficacy of accuracy and designed robustness metrics in distinguishing between robust and non-robust clusters. Modifications include retrieval-specific adjustments by excluding LM-induced non-robust examples.
4.2 Empirical Validation of GRAMMAR
@@
This section will validate design choices of identifying defective modules in the proposed framework.
The Importance of Modular Evaluation: Avoiding The Masking Effect of Comprehensive Assessments
Analysis presented in Table 2 underscores the value of segmenting the evaluation to specifically examine the retrieval process. The non-robust cluster actually receives higher scores (better performance) than the robust cluster regarding both the accuracy and robustness metrics. However, under the modular evaluation, the evaluation result correctly reflects the non-robust cluster. This underscores the importance of evaluating the retrieval function independently; failing to do so may allow language model inaccuracies to obscure the true performance metrics, potentially masking retrieval robustness issues. Although instances causing such distortions were not dominant in our dataset, they highlight a critical evaluation nuance.
The Importance of Excluding Gap Groups for Robustness Measurment
Highlighting the necessity of omitting gap groups, our ablation study implements two distinct approaches: 1) Comparing the Robustness Metric with Accuracy; 2) Comparative Analysis of Stratified Sampling and Unbalanced Sampling. Stratified sampling maintains a consistent $\lambda$ across both robust and non-robust clusters. By prompting the text template generator, an equal number of text queries are generated for each query logic. To explore the effects of unbalanced sampling, that can occur in the real world, we randomly reduce examples from query groups related to “clients” and “employees”.
Table 2 demonstrates that traditional accuracy metrics falter in distinguishing clusters based on vulnerability attributes, particularly when $\lambda$ varies between the hypothesis-driven cluster and its counterpart (Row 2 under the “Accuracy” category). Stratified sampling ameliorates this issue, enabling a precise identification of both robust and non-robust clusters (Row 4 under the “Accuracy” category). Importantly, the implementation of a specifically designed robustness metric ensures consistent and accurate identification of the robust cluster, irrespective of $\lambda$ variations (Row 2 and 4 under the “Robustness” category). This approach, integrating stratified sampling with our bespoke robustness metric, providing a reliable analysis.
@@ 5 Conclusions
In our work, we have developed an evaluation framework tailored for Retrieval-augmented Generation (RAG) systems within an industrial context, enhancing the reliability and interpretability of evaluations. Additionally, we have exposed the limitations of current reference-free evaluation methods, particularly their ineffectiveness in accurately assessing incorrect predictions.
@@
The entire evaluation framework, encompassing both the code and the prompts for the LLM, will be made open-source and available to the academic and industrial community. This initiative aims to facilitate the wider use and adaptation of our framework for research and real-world implementations. In terms of the systems and data evaluated to enable result replication, we plan to publish the model, including the documents intended for retrieval, as well as all templates and queries developed for the hypothetical Aurp scenario. However, the model designated for commercial use and its associated knowledge base will remain confidential and unpublished to honor non-disclosure agreements and protect proprietary information.
@@ 6 Limitations
The challenges related to the presence of multiple correct answers (refer to Appendix B.1) and the constraints in expressiveness posed by database schema and SQL (refer to Appendix B.4) ultimately result in difficulties in generating queries that demand multi-step reasoning and free-form responses. Nonetheless, as demonstrated in the case study, straightforward queries prove to be sufficient for identifying and debugging the intrinsic robustness issues within the models.
Appendix A Knowledge Structure and Representation
Appendix B Details of SQL Template Generation
B.1 Constraints/Guidelines for Creating SQL Templates
Selecting Attributes That Are Understandable to Humans
As shown in Table 3, the SQL template generator is required to follow the constraint: “The selected and condition columns in the query MUST BE MEANINGFUL and DESCRIPTIVE to ensure the queries are easily understood by non-technical users.”.
Avoiding Answer Multiplicity
B.2 LLM-Based, Human-Aided SQL Generation
We utilize GPT-4 as a SQL template generator for its state-of-the-art capabilities in SQL generation benchmarks (Liu et al., 2023). However, according to our results, LLMs could violate the constraints specified in the prompt. Therefore, the tool incorporates built-in functions that facilitate manual verification by saving, modifying, and reloading SQL templates. As detailed in Table 3, the profile prompt defines its purpose and establishes constraints for generating valid SQL templates.
The following process demonstrates the use of the developed tool to generate SQL templates:
1## a. Automatically generate SQL query templates
2sql_template_generator = SQLTemplateGenerator(connection_string, llm)
3sql_template_generator.set_system_msg()
4sql_templates = sql_template_generator.generate()
5sql_template_generator.save(sql_templates, root_dir, overwrite=True)
The output is saved in "SQLTemplateGenerator/sql_templates.txt" with the following SQL queries:
1SELECT Name FROM company WHERE Sales_in_Billion = ’[company.Sales_in_Billion]’;
2SELECT Industry FROM company WHERE Name = ’[company.Name]’;
3SELECT Headquarters FROM company WHERE Market_Value_in_Billion = ’[company.Market_Value_in_Billion]’;
4SELECT Nationality FROM people WHERE Age = ’[people.Age]’;
5SELECT Graduation_College FROM people WHERE Name = ’[people.Name]’;
6SELECT Year_working FROM employment WHERE Company_ID = ’[employment.Company_ID]’;
7SELECT People_ID FROM employment WHERE Year_working = ’[employment.Year_working]’;
To ensure the reliability of the generated SQL templates, they can be manually verified and modified, and new templates can be added. Modifications and reloading of templates can be performed using the following steps:
1# b. Manually modify SQL query templates (optional)
2
3# c. Load modified SQL query templatesn (optional)
4sql_templates = SQLTemplateGenerator.load(root_dir)
[…]
Below are the modified SQL queries as suggested in the §B.1:
1SELECT Industry FROM company WHERE Name = ’[company.Name]’;
2SELECT Headquarters FROM company WHERE Name = ’[company.Name]’;
3SELECT Nationality FROM people WHERE Name = ’[people.Name]’;
4SELECT Graduation_College FROM people WHERE Name = ’[people.Name]’;
5SELECT e.Year_working FROM employment e JOIN company c ON e.Company_ID = c.Company_ID JOIN people p ON e.People_ID = p.People_ID WHERE c.Name = ’[company.Name]’ AND p.Name = ’[people.Name]’;
B.3 Possibilities Of SQL Templates
One Table/Entity
Two Entities
B.4 Limitations
Scope of Using Automated SQL Generators
Queries Beyond SQL Expressiveness
Appendix C Breakdown of Scalable Data Generation
The proposed data generation process of generating queries via LLM, which includes both the creation of templates and the formulation of queries, can be succinctly represented by the formula (see Figure 2 for a visual representation):
Total Query $\text{Total Query Variations}=M\times N\times Q,$ (3)
Let’s break down each step of the proposed data generation process:
  • 1 schema $\Rightarrow$ M SQL Templates: The number of possible SQL templates (M) that can be generated from a defined schema is influenced by the SQL operators used, the number of tables, and the columns in the database schema. Each combination of tables and columns, along with different SQL operators (like SELECT, WHERE, JOIN), can lead to a unique SQL query template, e.g., the exponential complexity of the SELECT operator, and the combinatorial increase of possible predicates (See Appendix B.3 for details). Note that while the theoretical maximum is high, practical and meaningful queries will be a smaller subset, as specified in the case study and Appendix B.
Appendix D Semantic Group
Textual queries that originate from a specific SQL logic $q_{\text{sql}}$ are organized into a semantic group because of their underlying shared semantics. For instance, the SQL query “SELECT C.address FROM COMPANY C WHERE C.CompanyName = ’Carlton Innovation Precinct’“ results in the creation of a semantic group $Q=\{q_{t1},q_{t2},\ldots\}$, with each $q_{t}$ representing a distinct textual interpretation. Examples include $q_{t1}$ as “What is the address of ’Carlton Innovation Precinct’?” and $q_{t2}$ as “Where is ’Carlton Innovation Precinct’ located?”.
Appendix E Database Schemas Overview
(a) Aurecon.
(b) Aurp.
Figure 3: Entity-Relationship Diagrams for Data Generation
Attribute # Template # Row # Queries #
SQL SQL (Extra Constraints) Text SQL Text
Project 5 7 3 18 4 13 78
Employee 5 12 5 30 3 14 84
Client 2 6 2 12 4 6 36
Total / / / / / 33 198
(a) Aurecon
Attribute # Template # Row # Queries #
SQL Text SQL Text
Project 6 5 15 10 50 300
Employee 5 4 12 29 87 522
Client 3 2 6 10 20 120
Total / / / / 157 942
(b) Aurp
Table 4: Summary of DB Rows and Columns and Templates/Queries
Appendix F Synthetic Knowledge Base for Evaluating Retrieval-Augmented Language Models
The process of generating synthetic data for a synthetic environment involves creating comprehensive, detailed, and interconnected datasets for a fictitious company named Aurp, which specializes in structural engineering consultancy services. This multi-step procedure includes:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
This elaborate and interconnected data creation method is tailored for evaluating retrieval-augmented language models. It serves as a critical tool for exploring how these models handle data that is entirely fictional and thus not present in their training datasets. The synthetic nature of this environment provides a unique testbed for separating and assessing the distinct capabilities of the retrieval component and the language model component in isolation. Because the knowledge in the synthetic data does not appear in the training data of large language models.
Appendix G Assessing Scalability with Limited Entity Types, Rows, and Attributes
Our scalability assessment focused exclusively on the utilization of the SELECT operator and the WHERE clause for each table schema. As illustrated in Table 4(a) and 4(b), imposing constraints to generate relevant queries significantly reduces the potential quantity of SQL templates based on attribute count. Despite these restrictions, a minimal set of attributes and rows still yields hundreds of viable query-answer pairs.
Appendix H Prompt Templates for LLM-Based Evaluation

### v6

@@ 4.2 From Templates To Evaluation Data
Figure 2: Scaled Generation of Query-answer Pairs. Step 3 and step 4 are independent of each other and depend only on the SQL templates.
@@ 4.3 Modular Evaluation
Evaluating the overall performance of an RAG model can obscure the specific weaknesses or robustness of its constituent modules. This section introduces a modular evaluation protocol to address this.
Query Grouping and Tagging
The process involves grouping and tagging queries based on their semantics. Specifically, with GRAMMAR-Gen, textual queries $Q$ generated from a particular SQL query $q_{\mathrm{sql}}$ are organized into a semantic group. Depending on the model $M$’s performance, $q_{\mathrm{sql}}$ (and consequently, the group $Q$) are tagged into three principal categories, as shown in Figure 1.
@@
    $\forall q_{t}\in S,\neg M_{e}(M\left(q_{t}\right)),$
    where $M_{e}$ is the evaluation model, which returns “TRUE” if the model’s response $M\left(q_{t}\right)$ is correct.
@@
  • Non-robust Groups are characterized by model $M$’s ability to potentially answer the query semantics correctl, but with at least one query $q_{t}$ within $S$ that $M$ fails to predict correctly. This is formally denoted as:
    $\exists q_{t}\in S,M_{e}(M(q_{t}))\land\exists q_{t}\in S,\neg M_{e}(M(q_{t})).$
Assessing Retrieval Database via Gap Groups
Gap examples, which are incorrectly predicted instances within gap groups, highlight errors in the retrieval database that arise due to its limited coverage. The adequacy of the knowledge within the database can be quantified as follows:
$\text{Acc}_{\text{retrieval\_db}}=1-\frac{\text{Number of Gap Groups}}{\text{Total Number of Semantic Groups}}$ (1)
A low $\text{Acc}_{\text{retrieval\_db}}$ suggests the need for expanding the database.
Isolating Retrieval Database Errors
To ensure accurate evaluation of the retrieval and language models without interference from retrieval database errors (i.e., distinguishing knowledge gaps from robustness performance), two strategies have been developed:
  1. 1.
    Removing Gap Examples: To assess the other two modules, we calculate refined accuracy, denoted as $R$, by excluding gap examples:
@@
    $\begin{split}R=\frac{\text{\# with Correct Predictions}}{\text{Total \#}-\text{Total \# in Gap Groups}},\end{split}$ (2)
    where “#” represents “the number of instances”.
  2. 2.
    Balancing Gap Groups: To understand the advantage of this refined accuracy over baseline accuracy Acc, consider the following relationship:
    $\begin{split}\text{Acc}&=R\times\frac{\text{Total \#}-\text{Total \# in Gap Groups}}{\text{Total \#}}\\ &=R\times(1-\frac{\text{Total \# in Gap Groups}}{\text{Total \#}})\end{split}$ (3)
    This equation shows that a smaller normalized term, $\lambda=\frac{\text{Total \# in Gap Groups}}{\text{Total \#}}$, enhances evaluation accuracy. Intuitively, a smaller $\lambda$ implies fewer gap examples (errors propagated from the initial module), leading to more accurate results. Maintaining a consistent $\lambda$ across evaluation models or datasets ensures accurate assessments, such as for the hypothetically robust dataset $\mathcal{D}_{\text{robust}}$ and the non-robust dataset $\mathcal{D}_{\text{non-robust}}$ discussed in §5. With GRAMMAR-Gen, an equal number of text queries can be generated within each semantic group by adjusting the text template generator.
SQL Templates SQL Queries Text Templates Text Queries Text Queries (Balanced)
Short Long Short Long
Aurp 11 157 30 314 257 471 471
Spider-Open 5 57 10 397 436 570 570
Spider-Closed 5 57 10 426 430 570 570
Table 2: Statistics of Evaluation Datasets
Isolating Errors from Aurp Spider-Closed Spider-Open
Retrieval Database LLM Robust Non-Robust Robust Non-Robust Robust Non-Robust
Baseline: No Action No Action 0.27 0.51 0.26 0.28 0.67 0.87
Context Comparison 0.29 0.51 0.31 0.34 0.37 0.35
Remove Gap Groups No Action 0.95 0.96 0.58 0.44 0.98 0.96
Context Comparison 1 0.98 0.62 0.49 0.51 0.39
Balance Gap Examples No Action 0.27 0.27 0.28 0.21 0.69 0.68
Context Comparison 0.29 0.28 0.32 0.28 0.36 0.29
Table 3: Validating GRAMMAR’s tricks in in detecting non-robust retrieval module under a deliberately constructed RAG pipeline. Accuracy is calculated on selected examples across various types of semantic groups.
Identifying Non-Robust Retrieval or Non-Robust LM: A “Context Comparison” Approach
To determine whether robustness issues stem from the retrieval module or the language model, we focus on incorrectly predicted queries within the non-robust group, each referred to as $q_{\text{target}}$. Specifically, we pose the following question:
Does the retrieval module provide sufficient context for $q_{\text{target}}$?
By leveraging other semantically identical but correctly answered queries within the group, we can assess whether the context provided is sufficient. Specifically, we compare the document indices idxtarget associated with $q\text{target}$ to the indices of correctly answered queries, denoted as $I=\text{idx}1,\text{idx}2,\dots$. If idxtarget matches any index in $I$, the context provided by idxtarget is considered sufficient.
5 A Toy Case for Validation: Detecting A Vulnerable Retrieval
This section demonstrates the effectiveness of GRAMMAR on question answering tasks.
5.1 Experiment Setup
Evaluation Target
To validate our framework, we deliberately introduce a vulnerability in the retrieval component by using a basic keyword-matching retrieval approach. We use GPT3.5 for language generation. While this method is less common in practical applications, it is inherently weak in handling lengthy queries, making it an ideal candidate for testing our evaluation framework.
Evaluation Benchmarks and Datasets
Closed-domain RAG systems lack established benchmarks, especially because questions must be answered using documents containing knowledge that is not publicly available or known to public LLMs. To address this gap, we modify the Spider benchmark (Yu et al., 2018) and create two synthetic benchmarks with fictitious data. 1) Context-Augmented Spider (Open-Domain): We use the “company_employee” relational database from Spider to simulate a general real-world scenario. Documents for retrieval are synthesized using GPT-4 based on the database. 2) Context-Augmented Spider (Closed-Domain): This is similar to the first dataset, but with company names modified to fictitious entities that are either unused by specific companies or generic according to the latest GPT-4 models. 3) Aurp: To create a strictly closed-domain evaluation, we synthesize fictitious facts about a fictional company called Aurp. A relational database and corresponding retrieval documents are generated so that questions derived from the database can only be answered by the provided documents. Details on the data synthesis process are provided in Appendix G, and the database schemas are outlined in Appendix F. Using these three settings, datasets are generated by GRAMMAR-Gen to validate the design of GRAMMAR-Eval. The statistics of the generated datasets are summarized in Table 2.
5.2 Results
@@
This section validates the effectiveness of modular evaluation within the proposed framework in identifying defective modules.
GRAMMAR-Eval’s Effectiveness on Modular Evaluation
The results in Table 3 highlight the value of modular evaluation in identifying the defective retrieval. Regardless of the methods used to isolate errors from the retrieval database and LLMs, the accuracy consistently identifies the vulnerable retrieval module. The next three paragraphs will provide a more detailed analysis.
Preventing Forward Error Propagation
Preventing the propagation of knowledge deficits from retrieval databases is crucial. The two proposed strategies effectively address this issue. In all benchmarks, after either removing gap groups or balancing gap examples, accuracy on the robust text form consistently outperforms the non-robust text form (with the exception of Row 3 in Table 3, where the results are very close).
Impact of Backward Errors from LLMs on Retrieval Assessment
In some cases, such as Row 3 in Table 3, language model inaccuracies reduce the accuracy, masking the true performance of retrieval. While these instances are not prevalent in our dataset, they still distort the results. This distortion may be due to the stochastic nature of the LM’s performance, or, in a less favorable scenario, because the non-robust text forms used in retrieval inadvertently benefit the LM. Based on the results from the other two datasets, it is more likely the former.
Applicability of GRAMMAR to Open-Domain RAG
Table 3 also demonstrates the effectiveness of GRAMMAR for performing modular evaluation on the open-domain Spider dataset. However, in open-domain settings, the GPT model may generate accurate answers from its internal knowledge. This can lead to inaccuracies in assessing both the retrieval system’s knowledge gaps and the robustness of the retrieval mechanism. Specifically, the potential limitations in applying GRAMMAR-Eval to open-domain settings include: 1) Misjudgment of the Retrieval Database: Some gap examples in the retrieval database may go undetected, leading to an inflated Accretrieval_db. This issue is empirically demonstrated in Table 4, which compares Accretrieval_db before and after filtering out open-domain queries for Spider-Open. Closed-domain queries are identified by applying GRAMMAR to detect gap groups using answers from the LLM-only system. 2) Limitations of Context Comparison: Context comparison may fail when correctly answered examples do not actually receive sufficient context. This issue is evident in Table 3, where the accuracy decreased under “Context Comparison” compared to “No Action.” The decrease in accuracy is likely due to the introduction of incorrect context, which adds noise to the language model’s generation, thereby hindering its performance. In contrast, accuracy consistently improves for the two closed-domain benchmarks when context comparison is applied.
$\text{Acc}_{\text{retrieval\_db}}$
With Open-Domain Queries 0.53
Without Open-Domain Queries 0
Table 4: Wrong judgement on retrieval database with open-domain queries
@@ 6 Conclusions
In this work, we have introduced the GRAMMAR framework specifically designed for Retrieval-Augmented Generation (RAG) systems in closed-domain settings. The data generation module (GRAMMAR-Gen) improves the reliability of evaluation, while the evaluation protocol (GRAMMAR-Eval) enables detailed modular analysis.
@@
Our open-source Python package enables easy reproduction of all results, including the regeneration of experimental data and outcomes from our validation study. Detailed instructions on using the package are provided in Appendix H. Due to data privacy policies, the 198 examples and the RAG system from the sponsoring company are not included.
@@ 7 Limitations
The data generation process has certain limitations, such as its inability to handle scenarios with multiple correct answers (refer to Appendix D.1). Additionally, the expressiveness of the generated queries is constrained by the database schema and SQL (refer to Appendix D.3), limiting the ability to generate queries that require multi-step reasoning or free-form responses. However, as demonstrated in the case study, simple queries are sufficient for identifying and addressing intrinsic robustness issues within the models.
Besides, while the simple context comparison method performs well for closed-domain evaluation, it may need refinement for more complex retrieval tasks. This method was primarily used to validate the importance of modular evaluation and to establish a foundation, along with a reusable code base, for future development.
Appendix A Prompt Templates for LLM-Based Evaluation